# AutoScale: Automatic Prediction of Compute-optimal Data Composition for Training LLMs

### v2

@@ C.2 Scaling Latent Skills
$\mathbf{N^{(3)*}}=\mathbf{A}^{-1}\mathbf{A}\mathbf{N^{(2)*}}(\mathbf{N}^{(1)*})^{-1}\mathbf{N^{(2)*}}=\mathbf{N^{(2)*}}(\mathbf{N}^{(1)*})^{-1}\mathbf{N^{(2)*}}$
@@
The above result does not require identifying the latent skills or observing skill data compositions $\mathbf{K}$. Rather, the theorem gives that as long as the coefficient matrix $\mathbf{A}$ is invertible, the scaling of $\mathbf{N}$ complies to the same scaling law as in Sec. 3.3. ∎
@@
Appendix D Experimental Details and Additional Results
D.1 Experimental Details of Sec. 5.2
@@ Appendix D Experimental Details and Additional Results
GPT-2 Large is a variant of the GPT-2 architecture, featuring an embedding dimension of 1280, 36 transformer layers, and 20 attention heads. We rely on the Hugging Face Transformers library for implementation [Wolf et al., 2019]. Specific training hyperparameters are detailed in Table 2.
@@
Batch Size Per Device $1$
@@
D.2 Experimental Details of Sec. 5.3
@@
Mask Token Percentage $15$%
@@
D.3 Implementation Details of Baselines
@@
We followed the official implementation [^3] https://github.com/sangmichaelxie/doremi of DoReMi for our experiments. We evaluated two sets of reference domain weights: (1) the domain weights utilized in the LLaMA-2 paper Touvron et al. [2023] (referred to as LLaMA weights), and (2) uniform weights. Both the reference and proxy models have 120M parameters and are trained from scratch. We use GPT-2 tokenizer with a vocabulary size of roughly 50K. For LLaMA weights, we train each model for 20K, 50K and 200K steps for comparison. For uniform weights, we train each model for 10K, 20K and 50K steps. Refer to Table 4 for detailed hyperparameters. The effect of reference weights on the output DoReMi is discussed in Fig.9.
@@
Batch Size Per Device $8$
Devices $8$
@@ D.4 Evaluation Details
The following tasks are considered for downstream performance evaluation, in line with the setup from [Mehta et al., 2024, Gadre et al., 2024]. For few-shot tasks, the demonstrations are sampled at random.
  • BoolQ [Clark et al., 2019] consists of a question-answering format that requires binary yes/no answers.
  • HellaSwag [Zellers et al., 2019] challenges models on their ability to make commonsense inferences.
  • PIQA [Bisk et al., 2020] focuses on evaluating a model’s commonsense reasoning regarding physical interactions.
  • TruthfulQA [Lin et al., 2021] is designed to assess the ability of models to generate truthful and factual responses.
  • PubMedQA [Jin et al., 2019] offers a dataset for evaluating question-answering in the biomedical domain.
  • CrowsPairs-English [Nangia et al., 2020] tests models on their ability to identify and correct stereotypical biases in English text.
  • ARC-Easy [Clark et al., 2018] presents a set of relatively simpler scientific reasoning questions, aimed at evaluating a model’s basic understanding of scientific principles.
  • BigBench-Novel Concepts [Srivastava et al., 2022] serves as a test of the model’s creative abstraction skills, challenging it to make sense of scenarios that it could not have memorized during training.
@@
Batch Size Per Device $8$ or $300$
@@
Fig. 6 depicts AutoScale-predicted domain weights for training 774M Decoder-only LMs. Optimal data quantity for each domain grows in exponential-style functions with training data scale (left) where data sources with diverse samples (e.g., C4) are upweighted relative to domains with standard format (e.g., Wikipedia).
@@
Table 6 examines the domain-specific perplexity of GPT-2 Large trained on 3 billion tokens, respectively. Notably, AutoScale achieves the lowest average validation perplexity and significantly reduces the perplexity in the worst-performing domains.
@@
Domain/Method AutoScale DoReMi (Ref) Data Mixing LLaMA Uniform
Laws (ref) (30% more tokens)
Common Crawl 25.598 24.116 30.824 21.464 28.351
Github 7.482 6.678 5.845 7.376 5.784
Books 29.162 33.324 34.450 35.533 31.14
Wikipedia 18.828 17.154 26.795 21.110 19.57
C4 34.242 39.429 38.521 37.393 40.323
Stack Exchange 15.991 15.393 14.519 20.133 13.890
Arxiv 16.558 15.638 12.372 17.598 13.082
Average 21.123 21.676 23.333 22.944 21.736
Worst-domain 34.242 39.429 38.521 37.393 40.323
Table 6: Domain perplexity for 774M Decoder-only LMs trained for 3B tokens. AutoScale notably achieves the lowest average validation perplexity while also significantly decreasing worse-domain perplexity.
@@
Figure 8: Domain Weights used for training 774M Decoder-only LMs for 3B tokens. (Domain weights for Data Mixing Laws and DoReMi are from references [Ye et al., 2024] and [Fan et al., 2023], respectively, which are implemented on the same datasets/data domains with highly similar model architecture/model size/tokenizers.)
@@
D.6 Additional Results of Sec. 5.3
Fig. 10(b) shows the results on fitting validation loss with power-law functions, directly approximating how loss changes with each domain’s data quantity. Compared to BERT models trained with MLM (right), GPT models trained with CLM (left) demonstrate a much stronger response to domain reweighting. In final results, GPT/CLM achieved >2×>2$\times$> 2 × speed-up margins relative to uniform weights compared to BERT/MLM.
@@
Figure 10: Fitting validation loss with power-law functions, directly approximating how loss changes with each domain’s data quantity. Compared to BERT models trained with MLM (right), GPT models trained with CLM (left) demonstrate a much stronger response to domain reweighting. In final results, GPT/CLM achieved >2×>2$\times$> 2 × speed-up margins relative to uniform weights compared to BERT/MLM.
@@
D.7 Runtime Analysis
Training a GPT-2 large model from scratch for 3B tokens requires 15.5 hours on 8x NVIDIA A100 40GB SXM GPUs or 9 hours on 8x NVIDIA H100 80GB GPUs. Training time increases linearly with the number of training tokens on both types of GPUs.
Training BERT-base models takes 2 hours for every 18k steps on 4x NVIDIA A6000 48GB GPUs. Computational time grows linearly with the number of training steps.
Training reference models for DoReMi takes one hour for every 10K steps on 8x NVIDIA A6000 48GB GPUs. Computational time grows linearly with the number of training steps. Similar runtime for training proxy models for DoReMi.

### v5

@@ B.2 Scaling Latent Skills
$\mathbf{N^{*}}(N^{(3)})=\mathbf{A}^{-1}\mathbf{A}\mathbf{N^{*}}(N^{(2)})[(\mathbf{N}^{*}(N^{(1)}))^{-1}\mathbf{N^{*}}(N^{(2)})]^{k}=\mathbf{N^{*}}(N^{(2)})[(\mathbf{N}^{*}(N^{(1)}))^{-1}\mathbf{N^{*}}(N^{(2)})]^{k}$
@@
The above result does not require identifying the latent skills or observing skill data compositions $\mathbf{K}$. Rather, the theorem gives that as long as the coefficient matrix $\mathbf{A}$ is invertible, the scaling of $\mathbf{N}$ complies to the same scaling law as in Sec. 3.2.2. ∎
@@
Appendix C Experimental Details and Additional Results for Section 4, Evaluation
C.1 Experimental Details on GPT-2 Large (774M)
Evaluation We test the perplexity on the held-out dataset, comprising 10K samples each from the 7 domains. For downstream tasks, we include: BoolQ (Clark et al., 2019) (zero-shot), HellaSwag (Zellers et al., 2019) (zero-shot, 10-shot), PIQA (Bisk et al., 2020) (zero-shot), TruthfulQA (Lin et al., 2021) (zero-shot), PubMedQA (Jin et al., 2019) (10-shot), CrowsPairs (Nangia et al., 2020) (25-shot), and ARC-Easy (Clark et al., 2018) (zero-shot). Additionally, BBH Novel Concepts (Srivastava et al., 2022) task is added to the aggregated results for models trained beyond 10B tokens, making a total of 9 tasks. We select tasks that ensure the model’s performance surpasses random guessing, spanning from question answering and commonsense inference to bias identification and scientific problem solving. These tasks provide a comprehensive assessment of model performance (Mehta et al., 2024; Gadre et al., 2024). We adopt the evaluation framework from (Gao et al., 2021).
Baselines We report results for our methods (DDO and AutoScale ) and 6 baselines–Uniform, LLaMA weights (curated), DoReMi (LLaMA weights initialization), Data Mixing Laws from (Ye et al., 2024), DoReMi from Xie et al. (2024) (uniform initialization), and RegMix from Liu et al. (2024). Uniform weights uniformly sample data from all domains, resulting in the same number of training tokens from each domain. LLaMA weights are a set of curated domain weights heuristically tuned for training LLaMA-1/2 models. We implemented DoReMi proposed in (Xie et al., 2024). DoReMi trains two smaller-scale auxiliary models (proxy models). First, a reference model is trained with the dataset’s original domain weights, which are the LLaMA weights for RedPajama dataset. Then, optimized domain weights are obtained by using a proxy model to minimize the worst-case excess loss across different domains. We train both auxiliary models for 50K steps. Implementation details are available in App. C.3. Besides, we compare with 2 domain weights from existing literature, which are optimized on the same dataset, RedPajama, with similar Decoder-only LMs. Data Mixing Laws (Ye et al., 2024) first performs a grid search on the space of possible data mixtures and records evaluation loss for proxy models trained on these mixtures. Then, the loss is interpolated with exponential functions to find the optimal domain weights for the proxy model. DOGE (Fan et al., 2023) also implements DoReMi (Xie et al., 2024) with auxiliary models trained for 50K steps but with the reference model trained with uniform weights. RegMix (Liu et al., 2024) first trains an array of smaller, proxy models on different data mix and small data scales, abd fits a regression model between domain weights and evaluation loss. Then, the fitted regression model is used to predict the evaluation loss for all feasible domain weights to find the best-performing weights. We use the same pairs of domain weights and evaluation loss DDO used in optimizing domain weights for 774M Decoder-only LMs at 0.3B tokens to fit RegMix’s LightGBM regressor. The fitted LightGBM model is then used to optimize the evaluation loss over domain weights. We evaluate the model trained on these domain weights to present a complete landscape.
@@ C.1 Experimental Details on GPT-2 Large (774M)
GPT-2 Large is a variant of the GPT-2 architecture, featuring an embedding dimension of 1280, 36 transformer layers, and 20 attention heads. We rely on the Hugging Face Transformers library for implementation (Wolf et al., 2019). Specific training hyperparameters are detailed in Table 6.
@@
Batch Size Per Device 11
@@
C.2 Experimental Details on BERT (110M)
We evaluate the model’s MLM loss on held-out validation datasets, comprising 10K samples each from the 5 training domains. Additionally, as an auxiliary evaluation, we test the MLM loss on 3 non-training held-out domains. To be consistent with the perplexity loss used in CLM, we report the exponential cross-entropy loss for MLM. We evaluate the model’s task performance on GLUE benchmark (Wang et al., 2018) (with 8 diverse tasks for natural language understanding (NLU)) and SQuAD (Rajpurkar et al., 2016) (a large-scale QA dataset). Uniform weights are used as the baseline.
@@ C.2 Experimental Details on BERT (110M)
Mask Token Percentage 1515%
@@
C.3 Implementation Details for Baselines
@@
We followed the official implementation [^2] https://github.com/sangmichaelxie/doremi of DoReMi for our experiments. We evaluated two sets of reference domain weights: (1) the domain weights utilized in the LLaMA-2 paper Touvron et al. (2023) (referred to as LLaMA weights), and (2) uniform weights. Both the reference and proxy models have 120M parameters and are trained from scratch. We use GPT-2 tokenizer with a vocabulary size of roughly 50K. For LLaMA weights, we train each model for 20K, 50K and 200K steps for comparison. For uniform weights, we train each model for 10K, 20K and 50K steps. Refer to Table 8 for detailed hyperparameters. The effect of reference weights on the output DoReMi is discussed in Fig.9.
@@
Batch Size Per Device 88
Devices 88
@@ C.4 Evaluation Details
The following tasks are considered for downstream performance evaluation, in line with the setup from (Mehta et al., 2024; Gadre et al., 2024). For few-shot tasks, the demonstrations are sampled at random.
  • BoolQ (Clark et al., 2019) consists of a question-answering format that requires binary yes/no answers.
  • HellaSwag (Zellers et al., 2019) challenges models on their ability to make commonsense inferences.
  • PIQA (Bisk et al., 2020) focuses on evaluating a model’s commonsense reasoning regarding physical interactions.
  • TruthfulQA (Lin et al., 2021) is designed to assess the ability of models to generate truthful and factual responses.
  • PubMedQA (Jin et al., 2019) offers a dataset for evaluating question-answering in the biomedical domain.
  • CrowsPairs-English (Nangia et al., 2020) tests models on their ability to identify and correct stereotypical biases in English text.
  • ARC-Easy (Clark et al., 2018) presents a set of relatively simpler scientific reasoning questions, aimed at evaluating a model’s basic understanding of scientific principles.
  • BigBench-Novel Concepts (Srivastava et al., 2022) serves as a test of the model’s creative abstraction skills, challenging it to make sense of scenarios that it could not have memorized during training.
@@
Batch Size Per Device 88 or $300$
@@ C.5 Additional Results on GPT-2 Large (774M)
Fig. 6 depicts AutoScale-predicted domain weights for training 774M Decoder-only LMs (GPT-2 Large). Optimal data quantity for each domain grows in exponential-style functions with training data scale (left) where data sources with diverse samples (e.g., C4) are upweighted relative to domains with standard format (e.g., Wikipedia).
@@
Figure 8: Domain Weights used for training 774M Decoder-only LMs for 3B tokens. (Domain weights for Data Mixing Laws and DoReMi are from references (Ye et al., 2024) and (Fan et al., 2023), respectively, which are implemented on the same datasets/data domains with highly similar model architecture/model size/tokenizers.)
@@
C.6 Additional Results on BERT (110M)
Fig. 10(b) shows the results on fitting validation loss with power-law functions, directly approximating how loss changes with each domain’s data quantity. Compared to BERT models trained with MLM (right), GPT models trained with CLM (left) demonstrate a much stronger response to domain reweighting. In final results, GPT/CLM achieved $2\times$ speed-up margins relative to uniform weights compared to BERT/MLM.
@@ C.6 Additional Results on BERT (110M)
Figure 10: Fitting validation loss with power-law functions, directly approximating how loss changes with each domain’s data quantity. Compared to BERT models trained with MLM (right), GPT models trained with CLM (left) demonstrate a much stronger response to domain reweighting. In final results, GPT/CLM achieved $2\times$ speed-up margins relative to uniform weights compared to BERT/MLM.
@@
Appendix D Extended Discussions
Intuition for Stage 2 of AutoScale, Optimal mix projection.
Consider a stylized case where each training domain is independent. In this scenario, the validation performance attributable to each domain scales with the amount of training data from that same domain. Domains exhibit different scaling behaviors:
  • As the training budget increases, the loss associated with some domains (e.g., ’easier’ domains like standard texts from Wikipedia) decreases rapidly at first and then plateaus. If the total training data budget is small, it is sensible to allocate more budget to such domains, as they offer a greater initial reduction in validation loss.
  • Conversely, for other domains (e.g., diverse sources like Common Crawl), the loss decreases more slowly but steadily with an increasing training budget, continuing to provide benefits even when the loss from other domains has plateaued.
  • With larger compute budgets, it becomes more beneficial to allocate resources to these steadily improving domains rather than adding more data to ’easy’ domains that are already saturated. (As Figure 1 illustrates, optimal compositions are scale-dependent.) The key insight, which we formalize in a theorem, is that under certain independence assumptions, the evolution of optimal domain compositions with increasing training data budgets can be analytically derived from their individual scaling laws. This allows us to predict optimal compositions at larger data budgets.
We then relax the assumption of strict domain independence. We theoretically prove that if the validation data requires a number of independent ’latent skills,’ and data from each training domain contributes to one or more of these skills, our previous theory still holds in the same form. Thus, direct domain independence is not a strict requirement. This provides a complete theoretical grounding for our AutoScaletool, which captures the evolution of optimal domain weights with the training data budget and offers predictability for larger scales. Finally, our empirical evaluations demonstrate that the weights predicted by AutoScale offer clear advantages over domain weights optimized without considering this scale-dependence.the discussions and future work.
Number of predictors (scaling law components)
We conducted ablation studies with DML on optimizing domain weights under a fixed compute dudget, following the same experiment procedure in Section 4.1 to pre-train GPT-2 Large models on the RedPajama dataset. According to DML’s procedure, we fit a separate exponential function between domain weights and validation perplexity for each of the 7 domains, resulting in 7 separate predictors. When predicting for the average validation loss, we use each predictor to predict the validation perplexity for the respective domain and take the average of these predictions as the final output. When fitting on proxy training on random data mixtures, this prediction approach yields an average absolute relative error (AAR) = 4.46%. In comparison, similar to the treatment in the proposed DDO, we also fitted a single function between domain weights and the average validation perplexity using DML’s exponential formula. We experimented with using this function to directly predict the average validation perplexity. When fitting on proxy training on random data mixtures, this prediction approach yields an AAR = 1.61%.
When using the fitted predictor to derive optimal data mixtures, we realized DML’s formula does not allow optimize domain weights over a single predictor. For a validation domain, the predicted validation loss is given as the exponentiation of weighted average of domain weights plus some constant (Eq. (1) from Ye et al. (2024)). Since the domain weights are combined linearly before the exponentiation, if there is a single validation domain, minimizing the predicted loss will lead to a degenerate solution where all training data budgets is assigned to a single domain. Thus, it is necessary for DML to incorporate a number of separate predictors to introduce the crucial nonlinearity needed to model the interactions between different domains. The power-law formula in DDO is able to capture the nonlinearity in domain interactions in a single predictor, leading to lighter parametrization and enhanced efficiency